The length of a rectangle is 3x+4y feet, and the width is 2x-3y feet. compute the perimeter and the area.
Answer by fractalier(6550)

Here L = 3x + 4y and W = 2x - 3y.
Since Perimeter = 2L + 2W and Area = LW, we have
P = 2(3x + 4y) + 2(2x - 3y) =
P = 6x + 8y + 4x - 6y =
P = 10x + 2y
and
A = (3x + 4y)(2x - 3y) =
A = 6x^2 - xy - 12y^2 (via FOIL method)